# APCoTTA: Continual Test-Time Adaptation for Semantic Segmentation of Airborne LiDAR Point Clouds

### v1

@@ 2.4 Corruptions in 3D Semantic Segmentation
CTTA offers an effective solution to the challenge of continuously evolving domain shifts, but its performance relies on datasets that accurately capture real-world environmental variations. Ideally, collecting data under diverse weather conditions would provide the most realistic corruptions; however, this approach is costly and limited in scope. A more cost-effective alternative is synthesizing various corruption types on clean datasets. ImageNet-C [15] is the first to introduce various types of image distortions in image classification. Similarly, corrupted datasets have also emerged in fields such as image segmentation [18] and object detection [29]. For 3D point clouds, research has already established corruption benchmarks for tasks such as point cloud classification [35], object detection [9, 21], and semantic segmentation [21, 60]. However, they primarily focus on indoor or autonomous driving scenarios. Therefore, we similarly simulate corruption types on ALS point cloud datasets such as ISPRS and H3D, providing a benchmark for CTTA research in ALS point clouds.
Figure 1: The framework of our proposed method. At time step $t$, the current target domain batch $X_{t}$ is processed with both strong and weak augmentations and fed into the model $f_{\theta_{t}}$. First, the Dynamic Selection of Trainable Layers module identifies low-confidence layers as trainable while freezing the rest to mitigate catastrophic forgetting. Next, the Entropy-Based Consistency Loss module updates the trainable layers using high-confidence samples to reduce error accumulation. Finally, the Randomized Parameter Interpolation module randomly selects a subset of trainable parameters and blends them with their pretrained counterparts to further alleviate forgetting and enhance model stability.
@@ 3.1 Preliminaries and Overall Framework
Continual test‑time adaptation aims to adapt a source model $f_{\theta_{0}}$ pretrained on a labeled source domain $(\mathcal{X}^{S},\mathcal{Y}^{S})$ to a sequence of evolving, unlabeled target domains such as $\left[\mathcal{X}^{T}_{1},\mathcal{X}^{T}_{2},\cdots,\mathcal{X}^{T}_{n}\right]$. During this process, the source data remain inaccessible due to privacy and storage constraints, and the model has no knowledge of when domain shifts occur. Specifically, at each time step $t$, the current model $f_{\theta_{t}}$ receives unlabeled target‑domain inputs $x(t)$, makes predictions, and updates itself to $f_{\theta_{t+1}}$, which then processes the data $x(t+1)$ at time step $t+1$. At time step $t$, the complete model parameters $\theta_{t}$ consist of a set of parameter matrices, expressed as $\theta_{t}=\{\theta_{t}^{1},\theta_{t}^{2},\dots,\theta_{t}^{n},\dots,\theta_{t}^{N}\}$, where $\theta_{t}^{n}$ denotes the parameters of the $n$-th trainable layer at time step $t$, and $N$ denotes the total number of all trainable layers.
We propose APCoTTA, a continual test-time adaptation method for airborne LiDAR point cloud semantic segmentation. As illustrated in Fig. 1, APCoTTA consists of three modules: Dynamic Selection of Trainable Layers (DSTL), Entropy-Based Consistency Loss (EBCL), and Randomized Parameter Interpolation (RPI). We adopt a shared-weight framework where BatchNorm layers use the mean and variance of the current batch, and each input is augmented with both strong and weak perturbations. In the DSTL module, we compute the cross-entropy loss against a uniform distribution for each batch and update only the layers whose gradient norms fall below a predefined threshold, keeping the rest frozen. This strategy retains source domain knowledge and mitigates catastrophic forgetting. The EBCL module uses entropy as a confidence measure to filter out low-confidence samples and computes consistency loss only on high-confidence ones, reducing error accumulation. Finally, the RPI module randomly selects a subset of parameters from the trainable layers in each batch and interpolates them with the source domain parameters, balancing source and target knowledge to further alleviate forgetting.
@@ 3.2 Dynamic Selection of Trainable Layers
In CTTA, catastrophic forgetting refers to the loss of source domain knowledge caused by dynamic distribution shifts during adaptation to the evolving target domain, leading to overall performance degradation [53].
Recently, some methods achieve success in domain shifts by selectively adapting parts of the pretrained model, such as important parameters, layers, or network blocks [24, 33, 63, 28]. Inspired by them, we establish a confidence score to select confident and unconfident layers. We adaptively adjust the unconfident layers while freezing the confident layers to preserve source domain knowledge. Specifically, at any given inference time step t, we first compute the Kullback-Leibler (KL) divergence [22] between the model’s softmax predictions $\text{softmax}(f_{\theta_{t}}(x_{t}))$ and a uniform distribution u = [$\frac{1}{C}$, …, $\frac{1}{C}$] $\in$ $\mathbb{R}^{C}$, where C denotes the total number of classes in the dataset, and then perform backpropagation to obtain the gradients of each parameter.
$\mathcal{L}(\theta_{t})=\text{KL}(\text{{u}}||\text{softmax}(f_{\theta_{t}}(x_{t})))$ (1)
$\text{softmax}(f_{\theta_{t}}(x_{t}))=\frac{e^{f_{\theta_{t}}(x_{t})_{c}/T}}{\sum_{j=1}^{C}e^{f_{\theta_{t}}(x_{t})_{j}/T}}$
where $f_{\theta_{t}}(x_{t})$ represents the logits produced by the model at time step t, and T denotes the temperature scaling factor. Following [28], we set T to 50. The gradient with respect to the parameter $\theta_{t}$ is:
$\frac{\partial\mathcal{L}(\theta_{t})}{\partial\theta_{t}}=\frac{\partial\text{KL}(\text{{u}}||\text{softmax}(f_{\theta_{t}}(x_{t})))}{\partial\theta_{t}}=\frac{1}{C}\sum_{c=1}^{C}\frac{\partial\mathcal{L_{\text{CE}}}(f_{\theta_{t}}(x_{t}),c)}{\partial\theta_{t}}$ (2)
This shows that the gradient is equivalent to averaging the derivatives of the cross-entropy loss across all labels. Thus, we define the confidence score based on gradient norms:
$S(x_{t})=\lVert\frac{\partial\mathcal{L}(\theta_{t})}{\partial\theta_{t}}\rVert_{p}$ (3)
where $\lVert\cdot\rVert_{p}$ denotes $L_{p}$-norm. Following [16, 28], we adopt the L1 norm in this study. We set a confidence threshold $S_{0}$ and, during inference, update only the layers whose scores fall below the threshold, while keeping the other layers frozen.
@@ 3.3 Entropy-Based Consistency Loss
We input strong augmentation data $x_{t}^{s}$ and weak augmentation data $x_{t}^{w}$ separately into the same network $f_{\theta_{t}}$. Following [53, 28], we regularize the model by enforcing a consistency loss between the two predictions. However, some noisy samples may induce large gradients, potentially harming adaptation and leading to model collapse [32]. Given the success of [68, 52] in leveraging Shannon entropy [41] to select high-confidence samples, we also employ Shannon entropy to discard extremely low-confidence samples.
$H(p)=-\sum_{i=1}^{C}p_{i}\log p_{i}$ (4)
where $p_{i}$ is the probability of the $i$-th class (obtained from the softmax output of a model) and $C$ is the total number of classes. Unlike common practices that set thresholds for each class to select high-confidence samples, we set a threshold $\tau$ only to filter out extremely low-confidence samples and use the remaining samples for consistency loss.
$\mathcal{L}_{\text{con}}=\sum_{x_{t}}\mathbf{1}{\{H(p(x_{t}^{w}))<\tau\}}\cdot\left(-\sum_{i=1}^{C}p_{i}(x_{t}^{w})\log q_{i}(x_{t}^{s})\right)$ (5)
where $\tau$ is the entropy threshold, $\mathbf{1}{\{H(p(x_{t}^{w}))<\tau\}}$ is the indicator function, which is 1 when the entropy of $x_{t}^{w}$ is less than $\tau$, and 0 otherwise. By losing out low-confidence samples and retaining more accurate ones, we ensure gradient stability and mitigate error accumulation.
@@ 3.4 Randomized Parameter Interpolation
Through the DSTL and EBCL modules, we freeze most layers to retain source domain knowledge, mitigate catastrophic forgetting, and selectively backpropagate reliable predictions to reduce error accumulation. Since the target domain is unlabeled, prolonged adaptation may introduce errors and cause drift from the source domain. To further alleviate catastrophic forgetting, we randomly select a subset of parameters in the trainable layers selected by the DSTL module and apply weighted regularization toward the source domain weights.
@@
$\theta_{t}^{n}=M\odot\left(\alpha\theta_{0}^{n}+(1-\alpha)\theta_{t}^{n}\right)+(1-M)\odot\theta_{t}^{n},$ (7)
where $M$ is a mask tensor of the same shape as $\theta_{t}^{n}$, $p$ is the sampling probability for each element in $M$, $\alpha$ is the weighting coefficient that controls the contribution of the source domain weights, and $\odot$ denotes element-wise multiplication. Following [53], we set $p$ to 0.01. By weighting a subset of parameters in selected trainable layers with source domain parameters, we retain source knowledge while maintaining adaptability to the target domain, effectively avoiding deviation from the source and preventing overfitting to the target, thereby mitigating catastrophic forgetting.
@@ 4.1.1 Dataset Description
ISPRS dataset. The ISPRS Vaihingen 3D Semantic Labeling benchmark (ISPRS) dataset [37] is collected in the Stuttgart region of Germany using the Leica ALS50 system. The point cloud density ranges from 4 to 7 points/m² and includes nine categories: power lines, low vegetation, impervious surfaces, cars, fences, roofs, facades, shrubs, and trees. The dataset is divided into training and testing sets, containing 0.75 million and 0.41 million points, respectively, as shown in Fig. 2. The dataset comprises multiple scans with uneven point spacing across the region. To remove redundant points and ensure uniform point density, a subsampling grid size of d = 0.25 m is applied. The utilized features include X, Y, Z, Intensity.
H3D dataset. The Hessigheim 3D Benchmark (H3D) dataset [20] is collected using a RIEGL VUX1LR scanner mounted on a UAV, acquired through 11 longitudinal and several diagonal scan strips, with an average point density of approximately 800 pts/m². Higher densities occur in some areas due to additional scan strips. The data includes inherent LiDAR attributes such as XYZ coordinates, echo number, and reflectance, with point cloud colors derived via nearest-neighbor interpolation from a textured mesh. The training and validation sets contain approximately 59.4 million and 14.5 million points, respectively, as shown in Fig. 3, covering 11 categories: low vegetation, impervious surfaces, vehicles, urban furniture, roofs, facades, shrubs, trees, soil/gravel, vertical surfaces, and chimneys. Considering the high density of raw data, the data is downsampled to a 0.2 m resolution, utilizing features X, Y, Z, Reflectance.
@@ 4.1.2 Corruption Robustness Benchmarks
Following [9], we simulate corruptions at the weather and sensor levels, categorizing it into 7 curruption types grouped into five levels.
Weather-level corruptions: Strong Sunlight interferes with LiDAR signals, and we simulate this corruption by applying strong Gaussian noise to the point clouds. We do not consider extreme weather conditions such as rain, snow, or fog, as drones typically do not operate in such environments.
@@
By applying the aforementioned disturbance types to the ISPRS 3D test dataset and H3D val dataset, we establish two robustness benchmarks. Specifically, we apply the corruptions to their validation or test sets, as shown in Fig. 4,5, resulting in ISPRSC and H3DC. More detailed results are in the .1.

### v2

@@ 2.4 Corruptions in 3D Semantic Segmentation
CTTA offers an effective solution to the challenge of continuously evolving domain shifts, but its performance relies on datasets that accurately capture real-world environmental variations. Ideally, collecting data under diverse weather conditions would provide the highly realistic corruptions; however, this approach is costly and limited in scope. A more cost-effective alternative is synthesizing various corruption types on clean datasets. ImageNet-C [10] pioneers the introduction of various types of image distortions in image classification. Similarly, corrupted datasets have also emerged in fields such as image segmentation [13] and object detection [24]. For 3D point clouds, research has already established corruption benchmarks for tasks such as point cloud classification [27], object detection [6, 16], and semantic segmentation [16, 50]. However, they primarily focus on indoor or autonomous driving scenarios. To address this gap, we similarly simulate corruption types on ALS point cloud datasets such as ISPRS [29] and H3D [15], providing a benchmark for CTTA research in ALS point clouds.
Figure 1: Overview of the proposed framework. (a) The overall pipeline at time step $t$. The target domain batch $X_{t}$ is processed with both weak ($X_{t}^{w}$) and strong ($X_{t}^{s}$) augmentations and fed into the model $f_{\theta_{t}}$. The “Fire” and “Snowflake” icons explicitly indicate the trainable and frozen status of layers. (b) The Dynamic Selection of Trainable Layers (DSTL) module calculates the gradient norm based on the KL divergence to generate the selection mask $S_{l}$, identifying low-confidence layers for updates. (c) The Entropy-Based Consistency Loss (EBCL) module utilizes entropy statistics to filter and output high-confidence samples for reducing error accumulation. (d) The Randomized Parameter Interpolation (RPI) module randomly selects a subset of trainable parameters and blends them with their pretrained counterparts (Source Model) to further alleviate catastrophic forgetting.
@@ 3.1 Preliminaries and Overall Framework
Continual test-time adaptation aims to adapt a source model $f_{\theta_{0}}$ pretrained on a labeled source domain $(\mathcal{X}^{S},\mathcal{Y}^{S})$ with $C$ semantic categories to a sequence of evolving, unlabeled target domains such as $\{\mathcal{D}^{T}_{1},\mathcal{D}^{T}_{2},\dots,\mathcal{D}^{T}_{n}\}$. During this process, the source data remain inaccessible due to privacy and storage constraints, and the model has no knowledge of when domain shifts occur. Formally, at each time step $t$, the current model $f_{\theta_{t}}$ receives a batch of unlabeled target-domain inputs $X_{t}\in\mathbb{R}^{B\times N\times D_{in}}$ (where $B$ is the batch size, $N$ is the number of points, and $D_{in}$ denotes the input feature dimension), makes predictions, and updates itself to $f_{\theta_{t+1}}$. To enable robust adaptation, we adopt a teacher-student consistency strategy. Specifically, for each unlabeled target batch $X_{t}$, we generate two augmented views: a weakly augmented view denoted as $X_{t}^{w}$ (e.g., applying mild rigid transformations) and a strongly augmented view denoted as $X_{t}^{s}$ (e.g., applying aggressive jittering). At time step $t$, the complete model parameters $\theta_{t}$ consist of a set of parameter matrices, expressed as $\theta_{t}=\{\theta_{t}^{1},\theta_{t}^{2},\dots,\theta_{t}^{L}\}$, where $\theta_{t}^{l}$ denotes the parameters of the $l$-th trainable layer at time step $t$, and $L$ denotes the total number of all trainable layers. Unlike 2D images, ALS point clouds are characterized by irregular sparsity. Furthermore, severe distribution shifts arising from environmental variations during flight acquisition cause standard CTTA frameworks to suffer from severe local overfitting (e.g., to rural scenes), thereby inducing catastrophic forgetting.
We propose APCoTTA, a continual test-time adaptation framework for airborne LiDAR point cloud semantic segmentation. As illustrated in Fig. 1, APCoTTA consists of three modules: Dynamic Selection of Trainable Layers (DSTL), Entropy-Based Consistency Loss (EBCL), and Randomized Parameter Interpolation (RPI). We adopt a shared-weight framework where BatchNorm layers use the mean and variance of the current batch, and each input is augmented with both strong and weak perturbations. In the DSTL module, we compute the cross-entropy loss against a uniform distribution for each batch and update only the layers whose gradient norms fall below a predefined threshold, keeping the rest frozen. This strategy retains source domain knowledge and mitigates catastrophic forgetting. The EBCL module uses entropy as a confidence measure to filter out low-confidence samples and computes consistency loss only on high-confidence ones, reducing error accumulation. Finally, the RPI module randomly selects a subset of parameters from the trainable layers in each batch and interpolates them with the source domain parameters, balancing source and target knowledge to further alleviate forgetting.
@@ 3.2 Dynamic Selection of Trainable Layers
Compared with regular 2D images, ALS point clouds are irregular and sparse, characterized by heterogeneous local densities and geometric structures. As a result, feature statistics in point cloud networks are inherently less stable than those in image models, making adaptation more susceptible to catastrophic forgetting. This motivates the need for more robust parameter update strategies for point cloud networks during test-time adaptation. Through systematic exploration of parameter update strategies originally developed for image CTTA, we find that commonly used full-parameter update schemes are unsuitable for ALS point clouds, whereas gradient-based partial update methods [23] are more compatible with the point cloud networks. Based on this observation, we adopt and extend the gradient-driven layer selection mechanism to ALS tasks and develop the Dynamic Selection of Trainable Layers (DSTL) module, which automatically freezes layers that retain robust source domain priors while updating only those sensitive to geometric domain shifts.
Specifically, at any given inference time step t, we first compute the Kullback-Leibler (KL) divergence [17] between the model’s softmax predictions on the weakly augmented view, $\hat{y}_{t}^{w}=\text{softmax}(f_{\theta_{t}}(X_{t}^{w}))$, and a uniform distribution u = $\frac{1}{C}$, …, $\frac{1}{C}$] $\in$ $\mathbb{R}^{C}$, where C denotes the total number of classes in the dataset, and then perform backpropagation to obtain the gradients of each parameter.
$\mathcal{L}(\theta_{t})=\frac{1}{B}\sum_{i=1}^{B}\text{KL}(\text{{u}}||\hat{y}_{t,i}^{w})$ (1)
$\hat{y}_{t,i}^{w}=\frac{\exp({f_{\theta_{t}}(x_{t,i}^{w})_{c}/T)}}{\sum_{j=1}^{C}\exp({f_{\theta_{t}}(x_{t,i}^{w})_{j}/T)}}$
Here, $x_{t,i}^{w}$ is the $i$-th sample in the batch, $f_{\theta_{t}}(x_{t,i}^{w})$ represents the logits produced by the model at time step t, and $T$ is the temperature scaling factor (set to 50 following [23]). The gradient with respect to the parameter $\theta_{t}$ is:
$\frac{\partial\mathcal{L}(\theta_{t})}{\partial\theta_{t}}=\frac{\partial\text{KL}(\text{{u}}||\hat{y}_{t}^{w})}{\partial\theta_{t}}=\frac{1}{C}\sum_{c=1}^{C}\frac{\partial\mathcal{L}_{\text{CE}}(f_{\theta_{t}}(X_{t}^{w}),c)}{\partial\theta_{t}}$ (2)
This shows that the gradient is equivalent to averaging the derivatives of the cross-entropy loss across all labels. Thus, we then define the confidence score of the $l$-th layer based on its gradient norm:
$S_{l}(\theta_{t})=\frac{1}{K_{l}}\lVert\frac{\partial\mathcal{L}(\theta_{t})}{\partial\theta_{t}}\rVert_{p}$ (3)
where $\lVert\cdot\rVert_{p}$ denotes $L_{p}$-norm, and $K_{l}$ is the number of parameters in the $l$-th layer, used for normalization. Following [11, 23], we adopt the L1 norm in this study. In practice, to mitigate the bias caused by varying layer sizes, we normalize the gradient norm by the number of parameters in each layer. We set a confidence threshold $S_{0}$ and, during inference, update only the layers whose scores fall below the threshold, while keeping the other layers frozen.
@@ 3.3 Entropy-Based Consistency Loss
We input strong augmentation data $X_{t}^{s}$ and weak augmentation data $X_{t}^{w}$ separately into the same network $f_{\theta_{t}}$. Following [43, 23], we regularize the model by enforcing a consistency loss between the two predictions. However, error accumulation remains a significant risk due to the inherent physical noise in ALS data, such as solar background noise caused by intense sunlight, which leads to low-confidence predictions. Standard consistency losses indiscriminately align predictions across all samples [43], which can inadvertently propagate erroneous pseudo-labels from these low-confidence points. Given the success of [57, 42] in leveraging Shannon entropy [32] to select high-confidence samples, we introduce an entropy-based filtering mechanism to discard samples with extremely low confidence.
$H(\hat{y}_{t,i}^{w})=-\sum_{c=1}^{C}\hat{y}_{t,i,c}^{w}\log\hat{y}_{t,i,c}^{w}$ (4)
Unlike common practices that set thresholds for each class to select high-confidence samples, we employ a fixed global threshold $\tau$ to filter out extremely low-confidence samples. We define samples with entropy $H(\hat{y}_{t,i}^{w})\geq\tau$ as extremely low-confidence, representing high uncertainty where the model fails to produce a sharp probability distribution. Such samples often result from severe sensor noise or ambiguous boundaries common in ALS data. The choice of a fixed threshold over a dynamic one is crucial for stability. In ALS scenarios, batch statistics can fluctuate significantly due to local environmental factors (e.g., reflection saturation). A fixed $\tau$ provides a consistent standard, ensuring that the consistency loss is computed only on structurally reliable samples, thus mitigating unstable gradients caused by batch-level outliers. Finally, we use the remaining samples for consistency loss:
$\mathcal{L}_{\text{con}}=\frac{1}{N_{rel}}\sum_{i=1}^{B}\mathbb{I}(H(\hat{y}_{t,i}^{w})<\tau)\cdot\left(-\sum_{c=1}^{C}\hat{y}_{t,i,c}^{w}\log\hat{y}_{t,i,c}^{s}\right)$ (5)
where $\tau$ is the entropy threshold, $\mathbb{I}(\cdot)$ is the indicator function, and $N_{rel}$ is the number of reliable samples (to normalize the loss). By filtering out low-confidence samples and retaining more accurate ones, we ensure gradient stability and mitigate error accumulation.
@@ 3.4 Randomized Parameter Interpolation
Through the DSTL and EBCL modules, we freeze most layers to retain source domain knowledge, mitigate catastrophic forgetting, and selectively backpropagate reliable predictions to reduce error accumulation. Since the target domain is unlabeled, prolonged adaptation may introduce errors and cause drift from the source domain. While conventional methods utilize hard stochastic restoration by randomly resetting weights [43], such abrupt changes can disrupt the fine geometric calibration required for 3D point clouds. Consequently, we propose RPI as a soft regularization anchor. Instead of implementing hard resets, we randomly select a subset of parameters within the trainable layers identified by the DSTL module and apply weighted regularization to make them converge toward the source domain weights.
@@
$\theta_{t+1}^{l}=M_{l}\odot\left(\alpha\theta_{0}^{l}+(1-\alpha)\tilde{\theta}_{t}^{l}\right)+(1-M_{l})\odot\tilde{\theta}_{t}^{l}$ (7)
Here, $\tilde{\theta}_{t}^{l}$ denotes the parameters after the gradient update (via $\mathcal{L}_{\text{con}}$), and $\theta_{t+1}^{l}$ represents the final parameters for the next time step. $M_{l}$ is a binary mask sampled with probability $p$ (set to 0.01 [43]), $\alpha$ is a mixing coefficient, and $\odot$ denotes element-wise multiplication. By weighting a subset of parameters in selected trainable layers with source domain parameters, we retain source knowledge while maintaining adaptability to the target domain, effectively mitigating deviation from the source and mitigating overfitting to the target, thereby mitigating catastrophic forgetting.
@@ 4.1.1 Dataset Description
ISPRS dataset. The ISPRS Vaihingen 3D Semantic Labeling benchmark (ISPRS) dataset [29] is collected in the Stuttgart region of Germany using the Leica ALS50 system. The point cloud density ranges from 4 to 7 points/m² and includes nine categories: power lines, low vegetation, impervious surfaces, cars, fences, roofs, facades, shrubs, and trees. The dataset is explicitly divided into training and testing sets, containing 0.75 million and 0.41 million points, respectively (with an approximate split ratio of 65:35), as shown in Fig. 2. The dataset comprises multiple scans with uneven point spacing across the region. To remove redundant points and ensure uniform point density, a subsampling grid size of d = 0.25 m is applied. The utilized features include X, Y, Z, Intensity.
H3D dataset. The Hessigheim 3D Benchmark (H3D) dataset [15] is collected using a RIEGL VUX1LR scanner mounted on a UAV, acquired through 11 longitudinal and several diagonal scan strips, with an average point density of approximately 800 pts/m². Higher densities occur in some areas due to additional scan strips. The data includes inherent LiDAR attributes such as XYZ coordinates, echo number, and reflectance, with point cloud colors derived via nearest-neighbor interpolation from a textured mesh. The training and validation sets contain approximately 59.4 million and 14.5 million points, respectively (with an approximate split ratio of 80:20), as shown in Fig. 3, covering 11 categories: low vegetation, impervious surfaces, vehicles, urban furniture, roofs, facades, shrubs, trees, soil/gravel, vertical surfaces, and chimneys. Considering the high density of raw data, the data is downsampled to a 0.2 m resolution, utilizing features X, Y, Z, Reflectance.
Our selection of these two datasets is driven by their representativeness and complementarity, establishing a rigorous benchmark for CTTA. First, despite its release date, the ISPRS dataset remains a widely recognized “gold standard” for ALS segmentation and is actively used in related research, ensuring the continued relevance of our evaluation. Complementarily, the newer H3D (Hessigheim 3D) dataset is widely used in recent research, providing a modern and representative benchmark to further support our evaluation. Second, the two datasets represent fundamentally different real-world scenarios. ISPRS (Vaihingen) depicts a classic urban environment characterized by complex historic buildings and high-rise residential structures, acquired by manned aircraft with standard density (4–7 pts/m²). In sharp contrast, H3D captures a distinct village setting, acquired by modern UAV platforms with ultra-high density ($\approx$800 pts/m²). By evaluating on this diverse spectrum, ranging from legacy sparse urban data to modern dense rural data, we verify that APCoTTA is robust across significant variations in sensor platforms, point densities, and semantic environments, which is critical for reliable real-world deployment.
@@ 4.1.2 Corruption Robustness Benchmarks
Following [6], we simulate corruptions at the weather and sensor levels, categorizing it into 7 corruption types grouped into five levels.
Weather-level corruptions: Strong Sunlight interferes with LiDAR signals, and we simulate this corruption by applying strong Gaussian noise to the point clouds. We do not consider extreme weather conditions such as rain, snow, or fog, as UAVs typically do not operate in such environments.
@@
By applying the aforementioned disturbance types to the ISPRS 3D test dataset and H3D val dataset, we establish two robustness benchmarks. Specifically, we apply the corruptions to their validation or test sets, as shown in Fig. 4,5, resulting in ISPRSC and H3DC. More detailed results are provided in the .1.